OpenStudy (anonymous):

I am desperate i need help PLEASE!! Solve the following system of equations. 3x – 2y + z = –6 x – 3y + z = –3 2x + y – 2z = –7

OpenStudy (anonymous):

Can anyone help me??

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Do you know how to get the determinate of a matrix? If so, Cramer's rule is the easiest. x=Dx/D y=Dy/D z=Dz/D where D is thee determinate of the matrix Dx is the determinate of the matrix where the x values are substituted by the solutions, Dy is the determinate of the matrix where the y values are substituted by the solutions, and Dz is the determinate of the matrix where the z values are substituted by the solutions.

OpenStudy (phi):

are you set? stgreen posted the answer. elimination is how a computer would solve this. (Cramer's rule is computationally very expensive) But doing elimination by hand can be tedious and error prone, because you end up with fractions.

OpenStudy (phi):

Here is how I would do it x – 3y + z = –3 3x – 2y + z = –6 2x + y – 2z = –7 noticed I switched the order of the equations (that is allowed) so the one with a coeff on x is 1 (it makes things a little easier) write down the coeffs 1 -3 1 | -3 3 -2 1 | -6 2 1 -2 | -7

OpenStudy (anonymous):

i got the answe as -2,1,2?? is that not right??

OpenStudy (phi):

now "eliminate" the values below the 1 in the first column. you do this for the 2nd row by: 2nd row - 3* 1st row 3 -2 1 | -6 -3 - -9 -3 | --9 ______________________ 0 +7 -2 | +3 < -- this is your new 2nd row to eliminate the 1st entry in the bottom column: do 3rd row - 2*1st row 2 1 -2 | -7 -2 +6 -2 | +6 ___________________ 0 7 -4 | -1 <-- this is your new 3rd row so far we have 1 -3 1 | -3 0 +7 -2 | +3 0 7 -4 | -1 now use the 2nd row to eliminate the numbers below the 7 in the 3rd row you do this by 3rd row - 1*2nd row 0 7 -4 | -1 -0 -7 +2 | -3 _________________ 0 0 -2 | -4 <--- your new 3rd row you now have 1 -3 1 | -3 0 +7 -2 | +3 0 0 -2 | -4 if we put the letters back in, you can solve. start at the bottom and work up